Grade 11Physical Chemistry

Calculate frequency if wavenumber of radiation is 9000

Sai Dhanush Gedala
9 Years agoGrade 11

1 Answer

Nishant Kumar
9 Years ago
wavenumber= 1/wavelength;9000= 1/ wavelenghtwavelength= 1/9000 = 1.11*10^(-4)now, frequency= velocity of light/ wavelengthfrequency= 3*10^8/(1.11*10^-4) = 2.702* 10^12